Describe the transformation from the graph of f(x) = x + 3 to the graph of g(x) = x − 7.
Options:
The graph g(x) = x − 7 is the result of translating the graph of f(x) = x + 3 up 10 units.

The graph g(x) = x − 7 is the result of translating the graph of f(x) = x + 3 down 4 units.
The graph g(x) = x − 7 is the result of translating the graph of f(x) = x + 3 up 4 units.
OR
The graph g(x) = x − 7 is the result of translating the graph of f(x) = x + 3 down 10 units.

Answer by ewatrrr(24785)

f(x) -10 = g(x)
The graph g(x) = x − 7 is the result of translating the graph of f(x) = x + 3 down 10 units.